if cos theta=0.6,find sin theta and tan theta, by considering a right angled triangle with hypotonues of length 1,where theta is acute angle.
please can you shoe me the working out as well

Answer by Alan3354(69443)

if cos theta=0.6,find sin theta and tan theta, by considering a right angled triangle with hypotonues of length 1,where theta is acute angle.
--------------
sin^2 = 1 - cos^2
sin^2 = 1 - 0.36 = 0.64
sin = 0.8
------------
tan = sin/cos
tan = 4/3